the salary of two workers is in the ratio 3:4 they recieved 12,250 how much did each worker recieved
Answer by macston(5194)

Worker A is 3 parts and Worker B is 4 parts(3:4) for a total of 7 parts
Worker A= 3/7(100%)=42.9% of the total
Worker B= 4/7(100%)=57.1% of the total
ANSWER 1 Worker A received 42.9 % of 12250 or (0.429)(12250)=5255.25
ANSWER 2 Worker B received 57.1 % of 12250 or (0.571)(12250)=6994.75